EM022A

PI: Moscadelli
Title: YsOs/Water Masers

Session: Oct 1996 (session 3)
Frequency: 1.3 cm
Mode: C
Obs time (UT): 291 EU 16:00(17/10)-16:00(18/10)
Stations: EbJbMhNtOnSm
Processor: Bonn
EU Mk3 Tapes/St.: 3

Comment


To give feedback on this experiment, click here

Experiment Feedback

Metsahovi (Kirsi Karlamaa, Ari Mujunen, Kaj Wiik at Sat Oct 19 09:19:14 EDT 1996) Minor failures
It seems that drudg did not like this schedule since it emitted null-characters to the schedule file header and to some setup-procedure names. Affected scans were 3 - 6 (291 16:19 - 16:53). The phase cal signal was on till 291 17:10. Pointing computer crashed and was rebooted at 292 13:30 (affected data from about 292 13:26). Antenna cable wrap slew took from 292 10:35 - 10:37. Formatter time jumps corrected by switching off its fan and letting it run warmer...

Effelsberg (D Graham at Wed Oct 23 09:01:22 EDT 1996) Minor failures
Slewing times too optimistic: L1204G lost, NGC2071 behind hill

Noto (Corrado Trigilio at Wed Oct 30 05:13:45 EST 1996) Minor failures
Only some power fails and some late onsource. No other problems reported. Sky conditions: generally good.

Onsala (A. Polatidis, J. Conway at Tue Nov 5 05:16:52 EST 1996) Minor failures
Recording only LCP at this frequency, hence even VCs do not have any data. Problems with noise-cal switch during the middle of the epxperiment. Lost data day 292 UT 091100-091730 due to receiver problem. No other know problems.

Jodrell (Alastair Gunn at Thu Nov 7 06:11:18 EST 1996) Minor failures
Field system crashed and was restarted at 21:08UT, tape repositioned. Vacuum lost on recorder repeatedly from 12:10UT but data should be okay. Suspect this was due to old tape. IF attenuation needed changing frequently. Tapes MEDP1060 and MPIR1117 shipped to Bonn on 21/10/96 via Woodcroft. Weather cloudy throughout.

MPIfR (Walter Alef at Tue Sep 23 17:01:39 METDST 1997) Minor failures
Not observed by: Cambridge, Crimea, Medicina = 5 Stations left. About 1 hour lost at Jodrell Rest O.K.